# Medium separation scheme effects on the magnetized and cold two-flavor superconducting quark matter

### v1

@@ III Numerical Results
In order to discuss the numerical results, it is necessary to establish the model parameters, specifically the current quark mass $m_{c}$, the scalar coupling $G_{S}$, and the previously mentioned energy scale $\Lambda$. These parameters are determined by fitting them to empirical values: the pion decay constant $f_{\pi}=93.2$ MeV, the pion mass $m_{\pi}=135$ MeV, and the quark condensate in the vacuum, $\langle\bar{q}q\rangle^{1/3}_{0}=-250$ MeV. This procedure yields $G_{S}=4.75$ GeV-2, $m_{c}=4.99$ MeV, and $\Lambda=660$ MeV. For all the regularization schemes, the vacuum quark mass (which is also the mass scale for MSS) is $M_{0}\sim 302$ MeV, and the diquark coupling constant $G_{D}=0.75G_{S}$ is obtained via a Fierz transformation Buballa (2005). Additionally, a smoothness parameter $\alpha=0.01\Lambda$ is employed for the smooth functions of the Fermi-Dirac form, as in Mandal and Jaikumar (2017):
@@
We call this scheme nMFIR, following the notation of Avancini et al. (2019b).
@@ III.1 The finite magnetic field case
Fig. 1 compares the normalized effective quark mass $M/M_{0}$ as a function of the magnetic field at zero chemical potential (and, consequently, $\Delta=0$). It is well-established that the chiral condensate and the constituent quark mass, given their linear relation, increase with the magnetic field, a phenomenon known as magnetic catalysis Gusynin et al. (1996). As observed in this figure, both regularizations exhibit this phenomenon; however, the nMFIR scheme introduces strong unphysical oscillations that are more pronounced as the magnetic field increases, in contrast to the behavior obtained for MFIR Menezes et al. (2009). These artifacts are often mistaken for the well-known (vA-dH) oscillations, which are associated with discontinuities in the density, and are present at finite $\mu$, as can be seen in the following results.
@@
The analysis of the thermodynamic potential shows that the oscillations originate solely from the $\Omega_{1}$ term, which contains a theta function dependent on the magnetic field, thereby establishing an upper limit for the Landau-level summation. However, it notably vanishes at zero chemical potential. While $\Omega_{0}$ is independent of $eB$, the coupling of quark species to $\Delta$ in $\Omega_{\frac{1}{2}}$ eliminates the theta function in the $T=0$ limit. Consequently, for nonzero values of $\Delta$, the density is consistently non-zero across all Landau levels. This is the case for the $N_{c}=2$, where there are no theta functions in the zero temperature limit, and all the oscillations are strictly unphysical Duarte et al. (2016).
@@
Figure 2: Constituent quark mass $M$ (panel a) and diquark condensate $\Delta$ (panel b) as functions of the magnetic field, comparing nMFIR, MFIR and MFIR + MSS.
@@ III.2 The finite chemical potential case
This section presents the results obtained by varying the quark chemical potential at fixed magnetic field values. Since we are interested in the behavior of the physical quantities when the density increases, we compare the TRS and MSS results using MFIR in both cases. The panels in Fig. 3 illustrate the effective quark mass and the diquark condensate within the TRS (left panel) and MSS (right panel) frameworks at zero magnetic field. The effective quark mass $M$ exhibits qualitatively similar behavior across both methods. However, the diquark condensate reveals a significant discrepancy in the high chemical potential region.
Figure 3: Constituent quark mass $M$ and diquark condensate $\Delta$ as functions of the chemical potential, comparing TRS (panel a) and MSS (panel b) at $eB$ = 0.
@@
Fig. 4 illustrates the diquark condensate as a function of the chemical potential $\mu$ at finite magnetic field. We present only the $\Delta$ results, as the constituent quark mass exhibits the same qualitative behavior for both methods. This is consistent with Fig. 3 at $eB=0$, and this qualitative similarity persists for the finite values of $eB$ considered in this work. These results demonstrate that the behavior for $\Delta$ remains consistent with the $eB=0$ case: $\Delta$ increases monotonically with $\mu$ for MSS, while for TRS, it initially increases but eventually decreases to zero.
@@
This figure highlights the importance of the proper divergence treatment in the model when medium effects are included. The left panel shows that TRS predicts a normal phase with chiral symmetry restored and no color superconducting phase at large values of $\mu$. This prediction is an artifact of improper regularization, conflicting with other established approaches and with the expected physical picture at high chemical potential and low temperature. In this regime, the system is susceptible to the Cooper instability [^2] See Cooper (1956) for details on this instability and the Cooper theorem for Fermi systems. The analogous case for QCD was proposed as early as 1975 Collins and Perry (1975)., leading to the formation of Cooper pairs as the most stable state for deconfined quarks.
@@ III.3 Thermodynamic quantities
In this section, we present some thermodynamic quantities of the system, focusing on how the thermodynamics of the system is affected by the presence of an external magnetic field. An important outcome of this analysis is the emergence of anisotropy: in the presence of a strong magnetic field, the pressure splits into parallel and transverse components Ferrer et al. (2010); Ferrer and Hackebill (2022). The component of the pressure parallel to the direction of the magnetic field, $p^{\parallel}$, is obtained directly from the thermodynamic potential given in Eq. (LABEL:OmegaT0), in the same way as done at $B=0$:
@@
The figures in this section present the results for thermodynamic quantities for $eB=0.01$ and 0.1 GeV2, which correspond to $eB\approx 5\times 10^{17}$ G and $5\times 10^{18}$ G, respectively. The results obtained with TRS and MSS are compared, both implemented within the MFIR scheme. Panel (a) of Fig. 6 displays the magnetization, M, as a function of the chemical potential $\mu$. For the weaker field, both the TRS and MSS exhibit pronounced oscillations with increasing $\mu$, a behavior that can be associated with the filling of Landau levels. At $eB=0.10\penalty 10000\ \text{GeV}^{2}$ and nonvanishing diquark pairing, both regularization schemes converge toward similar behaviors, as Landau quantization dominates. At higher values of chemical potential, however, when $\Delta$ begins to decrease, their predictions differ: while TRS yields an eventually negative magnetization, the MSS scheme maintains a consistently positive magnetization across the entire range of $eB$ values considered, signaling a robust paramagnetic response of dense matter. The qualitative dependence of the magnetization on the chemical potential is also compatible with results obtained for cold magnetized quark matter within the nonlocal Nambu–Jona-Lasinio model Ferraris et al. (2025), suggesting that the main physical features remain stable across different effective descriptions. Although the diquark condensate vanishes within TRS at chemical potentials beyond the model scale, recent studies show that the MSS can be consistently extrapolated to $\mu>\Lambda$ in different contexts, yielding results in agreement with alternative approaches, including lattice simulations Lopes et al. (2025) and perturbative QCD calculations Pasqualotto et al. (2025); Azeredo et al. (2026). In Panel (b) of Fig. 6, one can see the equation of state $p_{N}^{\parallel}\times\varepsilon_{N}$. The EoS obtained within MSS is systematically softer than that produced by TRS, reflecting the proper separation of medium contributions without introducing artificial cutoffs. By preserving the physics near the Fermi surface, the MSS approach provides a more reliable description at high densities, with direct relevance for astrophysical applications, particularly the modeling of compact stars.
@@
Figure 6: Magnetization M as a function of chemical potential $\mu$ (panel a) and equation of state $p_{N}^{\parallel}\times\varepsilon_{N}$ (panel b) for a finite magnetic field.
@@
Panels (a) and (b) of Fig. 7 show the parallel and perpendicular components of the squared speed of sound, $c_{s}^{2}$, as functions of the baryon number density. The quantization of Landau levels generates van Alphen-de Haas oscillations, producing discontinuities in the density of states. This effect is more pronounced at lower magnetic fields, where more levels contribute, leading to higher-frequency oscillations. For $eB=0.10\penalty 10000\ \text{GeV}^{2}$, fewer levels contribute, and the behavior becomes smoother. Within MFIR, both TRS and MSS suppress spurious oscillations from regularization, while MSS preserves medium contributions near the Fermi surface, yielding physically consistent oscillations and a behavior for $c_{s}^{2}$ toward the conformal limit $1/3$. Furthermore, the qualitative dependence of the squared speed of sound on the baryon density agrees with recent nlNJL studies of cold magnetized quark matter Ferraris et al. (2025) and is also consistent with analyses that predict anisotropic sound velocities and a non-monotonic behavior approaching the conformal limit in strongly magnetized dense matter Ferrer and Hackebill (2023).
@@
Figure 7: The components of squared sound velocity $c_{s}^{2}$ as a function of the normalized baryon number density $n_{B}/n_{0}$, where $n_{0}$ denotes the nuclear saturation density. The figure shows the parallel component $(c_{s}^{\parallel})^{2}$ (panel a), while it shows the perpendicular component $(c_{s}^{\perp})^{2}$ (panel b). Results are shown for different regularization schemes in the presence of a finite magnetic field. The horizontal line indicates the limiting value $c_{s}^{2}=1/3$, characteristic of an ultrarelativistic gas.
@@ IV Discussion and Final Remarks
In the context of color superconducting phases under external magnetic fields, the application of MFIR is particularly important. The strong spurious oscillations observed in the order parameters when employing smooth functions dependent on $eB$ are frequently misinterpreted as the van Alphen–de Haas oscillations. While these genuine oscillations, associated with the filling of Landau levels at finite density, are present in the system, they are significantly less pronounced, as shown on solid and dashed curves of Fig. 2. The results obtained here clearly show that only the combination of MFIR and MSS fully removes the unphysical oscillations, leading to smooth and physically meaningful dependencies on both the magnetic field and density.
@@ Acknowledgements.
This work was partially supported by Conselho Nacional de Desenvolvimento Científico e Tecnológico (CNPq), Grants No. 312032/2023-4, 402963/2024-5 and 445182/2024-5 (R.L.S.F.); Fundação de Amparo à Pesquisa do Estado do Rio Grande do Sul (FAPERGS), Grants Nos. 24/2551-0001285-0 (R.L.S.F.) and 23/2551-0000791-6 and 23/2551-0001591-9 (D.C.D.); Coordenação de Aperfeiçoamento de Pessoal de Nível Superior - Brasil (CAPES) - Finance Code 001 (F.X.A.). The work is also part of the project Instituto Nacional de Ciência e Tecnologia - Física Nuclear e Aplicações (INCT - FNA), Grants No. 464898/2014-5 and 408419/2024-5, and supported by the Serrapilheira Institute (grant number Serra - 2211-42230). R. L. S. F. acknowledges the kind hospitality of the Center for Nuclear Research at Kent State University, where part of this work was done.

### v2

@@ III Numerical Results
In order to discuss the numerical results, it is necessary to establish the model parameters, specifically the current quark mass $m_{c}$, the scalar coupling $G_{S}$, and the previously mentioned energy scale $\Lambda$. These parameters are determined by fitting them to empirical values; the pion decay constant $f_{\pi}=93.2$ MeV, the pion mass $m_{\pi}=135$ MeV, and the quark condensate in the vacuum, $\langle\bar{q}q\rangle^{1/3}_{0}=-250$ MeV. This procedure yields $G_{S}=4.75$ GeV-2, $m_{c}=4.99$ MeV, and $\Lambda=660$ MeV. For all the regularization schemes, the vacuum quark mass (which is also the mass scale for MSS) is $M_{0}\sim 302$ MeV, and the diquark coupling constant $G_{D}=0.75G_{S}$ is obtained via a Fierz transformation Buballa (2005). Additionally, a smoothness parameter $\alpha=0.01\Lambda$ is employed for the smooth functions of the Fermi-Dirac form, as in Mandal and Jaikumar (2017):
@@
We call this scheme nonmagnetic field independent regularization (nMFIR), following the notation of Avancini et al. (2019b).
@@ III.1 The finite magnetic field case
Figure 1 compares the normalized effective quark mass $M/M_{0}$ as a function of the magnetic field at zero chemical potential (and, consequently, $\Delta=0$). It is well-established that the chiral condensate and the constituent quark mass, given their linear relation, increase with the magnetic field, a phenomenon known as magnetic catalysis Gusynin et al. (1996). As observed in this figure, both regularizations exhibit this phenomenon; however, the nMFIR scheme introduces strong unphysical oscillations that are more pronounced as the magnetic field increases, in contrast to the behavior obtained for MFIR Menezes et al. (2009). These artifacts are often mistaken for the well-known (vA-dH) oscillations, which are associated with discontinuities in the density, and are present at finite $\mu$, as can be seen in the following results.
@@
The analysis of the thermodynamic potential shows that the oscillations originate solely from the $\Omega_{1}$ term, which contains a theta function dependent on the magnetic field, thereby establishing an upper limit for the Landau-level summation. However, it notably vanishes at zero chemical potential. While $\Omega_{0}$ is independent of $eB$, the coupling of quark species to $\Delta$ in $\Omega_{\frac{1}{2}}$ eliminates the theta function in the $T=0$ limit. Consequently, for nonzero values of $\Delta$, the density is consistently nonzero across all Landau levels. This is the case for the $N_{c}=2$, where there are no theta functions in the zero temperature limit, and all the oscillations are strictly unphysical Duarte et al. (2016).
@@
Figure 2: Constituent quark mass $M$ [panel (a)] and diquark condensate $\Delta$ [panel (b)] as functions of the magnetic field, comparing nMFIR, MFIR, and MFIR + MSS.
@@ III.2 The finite chemical potential case
This section presents the results obtained by varying the quark chemical potential at fixed magnetic field values. Since we are interested in the behavior of the physical quantities when the density increases, we compare the TRS and MSS results using MFIR in both cases. The panels in Fig. 3 illustrate the effective quark mass and the diquark condensate within the TRS [panel (a)] and MSS [panel (b)] frameworks at zero magnetic field. The effective quark mass $M$ exhibits qualitatively similar behavior across both methods. However, the diquark condensate reveals a significant discrepancy in the high chemical potential region.
Figure 3: Constituent quark mass $M$ and diquark condensate $\Delta$ as functions of the chemical potential, comparing TRS [panel (a)] and MSS [panel (b)] at $eB$ = 0.
@@
Figure 4 illustrates the diquark condensate as a function of the chemical potential $\mu$ at finite magnetic field. We present only the $\Delta$ results, as the constituent quark mass exhibits the same qualitative behavior for both methods. This is consistent with Fig. 3 at $eB=0$, and this qualitative similarity persists for the finite values of $eB$ considered in this work. These results demonstrate that the behavior for $\Delta$ remains consistent with the $eB=0$ case; $\Delta$ increases monotonically with $\mu$ for MSS, while for TRS, it initially increases but eventually decreases to zero.
@@
This figure highlights the importance of the proper divergence treatment in the model when medium effects are included. The left panel shows that TRS predicts a normal phase with chiral symmetry restored and no color superconducting phase at large values of $\mu$. This prediction is an artifact of improper regularization, conflicting with other established approaches and with the expected physical picture at high chemical potential and low temperature. In this regime, the system is susceptible to the Cooper instability, [^2] See Cooper (1956) for details on this instability and the Cooper theorem for Fermi systems. The analogous case for QCD was proposed as early as 1975 Collins and Perry (1975). leading to the formation of Cooper pairs as the most stable state for deconfined quarks.
@@ III.3 Thermodynamic quantities
In this section, we present some thermodynamic quantities of the system, focusing on how the thermodynamics of the system is affected by the presence of an external magnetic field. An important outcome of this analysis is the emergence of anisotropy; in the presence of a strong magnetic field, the pressure splits into parallel and transverse components Ferrer et al. (2010); Ferrer and Hackebill (2022). The component of the pressure parallel to the direction of the magnetic field, $p^{\parallel}$, is obtained directly from the thermodynamic potential given in Eq. (LABEL:OmegaT0), in the same way as done at $B=0$,
@@
The figures in this section present the results for thermodynamic quantities for $eB=0.01$ and 0.1 GeV2, which correspond to $eB\approx 5\times 10^{17}$ G and $5\times 10^{18}$ G, respectively. The results obtained with TRS and MSS are compared, both implemented within the MFIR scheme. Panel (a) of Fig. 6 displays the magnetization, M, as a function of the chemical potential $\mu$. For the weaker field, both the TRS and MSS exhibit pronounced oscillations with increasing $\mu$, a behavior that can be associated with the filling of Landau levels. At $eB=0.10\penalty 10000\ \text{GeV}^{2}$ and nonvanishing diquark pairing, both regularization schemes converge toward similar behaviors, as Landau quantization dominates. At higher values of chemical potential, however, when $\Delta$ begins to decrease, their predictions differ: while TRS yields an eventually negative magnetization, the MSS scheme maintains a consistently positive magnetization across the entire range of $eB$ values considered, signaling a robust paramagnetic response of dense matter. The qualitative dependence of the magnetization on the chemical potential is also compatible with results obtained for cold magnetized quark matter within the nonlocal Nambu–Jona-Lasinio model Ferraris et al. (2025), suggesting that the main physical features remain stable across different effective descriptions. Although the diquark condensate vanishes within TRS at chemical potentials beyond the model scale, recent studies show that the MSS can be consistently extrapolated to $\mu>\Lambda$ in different contexts, yielding results in agreement with alternative approaches, including lattice simulations Lopes et al. (2025) and perturbative QCD calculations Pasqualotto et al. (2025); Azeredo et al. (2026). In panel (b) of Fig. 6, one can see the equation of state $p_{N}^{\parallel}\times\varepsilon_{N}$. The equation of state obtained within MSS is systematically softer than that produced by TRS, reflecting the proper separation of medium contributions without introducing artificial cutoffs. By preserving the physics near the Fermi surface, the MSS approach provides a more reliable description at high densities, with direct relevance for astrophysical applications, particularly the modeling of compact stars.
@@
Figure 6: Magnetization M as a function of chemical potential $\mu$ [panel (a)] and equation of state $p_{N}^{\parallel}\times\varepsilon_{N}$ [panel (b)] for a finite magnetic field.
@@
Panels (a) and (b) of Fig. 7 show the parallel and perpendicular components of the squared speed of sound, $c_{s}^{2}$, as functions of the baryon number density. The quantization of Landau levels generates van Alphen-de Haas oscillations, producing discontinuities in the density of states. This effect is more pronounced at lower magnetic fields, where more levels contribute, leading to higher-frequency oscillations. For $eB=0.10\penalty 10000\ \text{GeV}^{2}$, fewer levels contribute, and the behavior becomes smoother. Within MFIR, both TRS and MSS suppress spurious oscillations from regularization, while MSS preserves medium contributions near the Fermi surface, yielding physically consistent oscillations and a behavior for $c_{s}^{2}$ toward the conformal limit $1/3$. Furthermore, the qualitative dependence of the squared speed of sound on the baryon density agrees with recent nonlocal Nambu–Jona-Lasinio (nlNJL) studies of cold magnetized quark matter Ferraris et al. (2025) and is also consistent with analyses that predict anisotropic sound velocities and a nonmonotonic behavior approaching the conformal limit in strongly magnetized dense matter Ferrer and Hackebill (2023).
@@
Figure 7: The components of squared sound velocity $c_{s}^{2}$ as a function of the normalized baryon number density $n_{B}/n_{0}$, where $n_{0}$ denotes the nuclear saturation density. The figure shows the parallel component $(c_{s}^{\parallel})^{2}$ [panel (a)], while it shows the perpendicular component $(c_{s}^{\perp})^{2}$ [panel (b)]. Results are shown for different regularization schemes in the presence of a finite magnetic field. The horizontal line indicates the limiting value $c_{s}^{2}=1/3$, characteristic of an ultrarelativistic gas.
@@ IV Discussion and Final Remarks
In the context of color superconducting phases under external magnetic fields, the application of MFIR is particularly important. The strong spurious oscillations observed in the order parameters when employing smooth functions dependent on $eB$ are frequently misinterpreted as the van Alphen-de Haas oscillations. While these genuine oscillations, associated with the filling of Landau levels at finite density, are present in the system, they are significantly less pronounced, as shown on solid and dashed curves of Fig. 2. The results obtained here clearly show that only the combination of MFIR and MSS fully removes the unphysical oscillations, leading to smooth and physically meaningful dependencies on both the magnetic field and density.
@@ Acknowledgements.
This work was partially supported by Conselho Nacional de Desenvolvimento Científico e Tecnológico (CNPq), Grants No. 312032/2023-4, No. 402963/2024-5 and 445182/2024-5 (R.L.S.F.); Fundação de Amparo à Pesquisa do Estado do Rio Grande do Sul (FAPERGS), Grants No. 24/2551-0001285-0 (R.L.S.F.), No. 23/2551-0000791-6 and No. 23/2551-0001591-9 (D.C.D.); Coordenação de Aperfeiçoamento de Pessoal de Nível Superior - Brasil (CAPES) - Finance Code 001 (F.X.A.). The work is also part of the project Instituto Nacional de Ciência e Tecnologia - Física Nuclear e Aplicações (INCT - FNA), Grants No. 464898/2014-5 and No. 408419/2024-5, and supported by the Serrapilheira Institute (Grant No. Serra - 2211-42230). R. L. S. F. acknowledges the kind hospitality of the Center for Nuclear Research at Kent State University, where part of this work was done.
Data Availability
The data that support the findings of this article are not publicly available. The data are available from the authors upon reasonable request.